# SAIF: Sparse Adversarial and Imperceptible Attack Framework

### v3

@@ 7 Ablation Studies
Attack Type Dataset ASR $\uparrow$ $||\mathbf{p}||_{0}/m$downarrow
@@ 8 Conclusion
Ethical Statement
Adversarial attacks expose the fragility of DNNs. Our work aims at demonstrating that a highly imperceptible adversarial attack can be generated for natural images. This provides a new benchmark for the research community to test the robustness of the learning algorithms. A straightforward defense strategy can be using the adversarial examples generated by SAIF for adversarial training. We leave more advanced solutions for future exploration.
@@ C.1 Dataset and model
We test SAIF and the existing sparse attack algorithms on the CIFAR-10 dataset (Krizhevsky et al., 2009). The dataset comprises $[32\times 32]$ RGB images belonging to 10 classes. We evaluate all algorithms on 10,000 samples from the test set.
@@ Appendix D Visual interpretability of SAIF
Table 9: Quantitative evaluation of untargeted attack on CIFAR-10. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation ‘$\epsilon$’.
@@
Table 10: Quantitative evaluation of targeted attack on CIFAR-10. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation ‘$\epsilon$’.

### v4

@@ 7 Ablation Studies
Attack Type Dataset ASR $\uparrow$ $\mathbf{p}||_{0}/m \downarrow$
@@ 8 Conclusion
Discussion and Ethics Statement
Adversarial attacks expose the fragility of DNNs. Our work aims to demonstrate that a highly imperceptible adversarial attack can be generated for natural images. This provides a new benchmark for the research community to test the robustness of the learning algorithms. A straightforward defense strategy can be using the adversarial examples generated by SAIF for adversarial training. We leave more advanced solutions for future exploration.
Moreover, in principle, SAIF can be extended to other input modalities, with appropriate adaptations for the target domain. For example, extending SAIF to video data would require adding a temporal consistency constraint, alongside the existing sparsity and magnitude constraints, to preserve the imperceptibility of perturbations. We leave a detailed exploration of these potential extensions to future work.
Acknowledgments
This project was supported by NIH/NCI grant R01CA240771.
@@ C.1 Dataset and model
We test SAIF and the existing sparse attack algorithms on the CIFAR-10 dataset (Krizhevsky et al., 2009). The dataset comprises $32\times 32]$ RGB images belonging to 10 classes. We evaluate all algorithms on 10,000 samples from the test set.
@@ Appendix D Visual interpretability of SAIF
Table 9: Quantitative evaluation of untargeted attack on CIFAR-10. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation $\epsilon$’.
@@
Table 10: Quantitative evaluation of targeted attack on CIFAR-10. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation $\epsilon$’.